# MOSS: Efficient and Accurate FP8 LLM Training with Microscaling and Automatic Scaling

### v1

@@ 4.2 LLM Pre-Training
MOSS 241.8(+47.5%) 52.8 64.7 59.4
We begin by comparing the performance of OLMo-7B models (Groeneveld et al., 2024) pre-trained using FP8 mixed precision with those trained using BF16. The pre-training loss over training steps is shown in Figure 5. All models use identical training configurations and hyperparameters as specified in the official report. As illustrated in Figure 5, the loss curves of BF16, COAT, and our proposed method, MOSS, closely align, exhibiting nearly identical trends. These results clearly demonstrate that the proposed FP8 mixed-precision scheme can match the training performance of the widely adopted higher-precision BF16 baseline.
4.3 LLM Fine-Tuning
We further evaluate the proposed FP8 low-precision training scheme in the context of instruction tuning for LLMs. To ensure a fair comparison, we adopt the same fine-tuning setup as ToolBench (Guo et al., 2024), using the open-source LLaMA-2-7B (Grattafiori et al., 2024) as the base model. Experiments are conducted on the math-focused MAmmoTH dataset (Yue et al., 2023), fine-tuned for 5 epochs. During fine-tuning, MOSS matches the performance of the BF16 baseline, with loss curves that are either closely aligned or exhibit slightly improved stability (Figure 6). Downstream task accuracies are reported in Figure 6. These results further demonstrate the numerical precision and robustness of MOSS, confirming that the proposed FP8 quantization scheme generalizes well beyond pre-training and remains effective in fine-tuning settings.
@@ 4.4 Ablation Study
To evaluate the efficiency of the GEMM kernel in MOSS, we conduct an ablation study comparing our implementation with several state-of-the-art FP8 quantization schemes, including Transformer Engine (TE)(Nvidia, 2025), DeepGEMM from DeepSeek(Zhao et al., 2025), and COAT (Xi et al., 2024). This evaluation focuses on both raw GEMM performance and the practical trade-offs in quantization and hardware deployment.
Table 4: Runtime of quantized FP8 GEMM on regular GPUs.
@@
As shown in Table 4, MOSS achieves GEMM performance that is comparable to TE. Importantly, MOSS significantly outperforms COAT in GEMM throughput, highlighting the efficiency of our low-overhead quantization strategy. In comparison to DeepGEMM, our GEMM kernel runtime is relatively higher. This is expected, as DeepSeek leverages the Increasing Accumulation Precision technique to reduce dequantization overhead and employs extensive hardware-specific optimizations. In contrast, our implementation is not tailored to Hopper GPUs, which lack the native Tensor Core support for MX data format. Despite this, MOSS remains competitive in performance across widely available hardware and maintains robust generalizability without relying on hardware-specific enhancements.
@@ 5.1 Low-Precision Training
Low-precision training refers to using low-precision bits, such as FP16/BF16, in the training process, as opposed to using FP32 full-precision throughout the training process. It is one of the most prominent directions in modern deep learning, as it offers reductions in computation costs, memory footprint and communication overhead (Micikevicius et al., 2017). Most existing training frameworks, e.g., Megatron-LLM (Shoeybi et al., 2019) and DeepSpeed (Rasley et al., 2020), adopt BF16 in LLM training by default. With the release of Nvidia Hopper GPUs, FP8 training (Micikevicius et al., 2022) becomes the next-generation low-precision technique. Transformer Engine (Nvidia, 2025) from NVIDIA is the first framework designed for FP8 mixed-precision training, however, it only supports FP8 compute for linear layers while leaving all other operations in full precision. FP8-LM (Peng et al., 2023) extends FP8 quantization to gradients and first-order optimizer states, further improving the training throughput. However, they fail to reduce the memory footprint of activations, not fully exploiting the potential of memory and communication benefits of FP8 in fully sharded distributed training scenarios. COAT (Xi et al., 2024) and DeepSeek-V3 (Liu et al., 2024) both extend the per-group FP8 quantization to activations, achieving mixed granularity quantization. However, the fine-grained quantization strategy incurs additional computation cost in the dequantization process.
@@ 5.2 Scaling Techniques
Quantization transforms high-precision values within a given range into lower-precision values of a different range by tensor scaling. Scaling factor determination during training typically follows one of two strategies: Delayed scaling estimates the current scaling factor using a historical record of maximum absolute values from previous iterations. This approach assumes statistical consistency across iterations, making it vulnerable to outliers that violate this assumption and potentially destabilize training (Fishman et al., 2024). In contrast, just-in-time scaling dynamically adjusts scaling factors based on the current data distribution, typically using on-the-fly max-reduction. While more adaptive, this method introduces significant computational overhead—particularly in FP8 training—where multiple memory accesses can offset the performance gains of low-precision computation. Beyond these non-automatic scaling approaches, Unit Scaling (Blake et al., 2023) recently introduced an automatic scaling method that maintains unit variance in weights and activations by applying static scaling constants throughout the network. However, Unit Scaling requires manual modification of the computational graph to insert unit-scaled operations and preserves certain “critical matmuls” (e.g., attention output projection and FFN down projection) in BF16 for stability. In contrast, MOSS offers a generic and automated scaling mechanism that integrates seamlessly into modern LLMs, achieving full FP8 precision without requiring precision fallbacks or manual graph modifications.
@@ Appendix C Discussion of Bound in AdamW Optimizer
Compared to Adam optimizer (Kingma & Ba, 2014), AdamW (Loshchilov & Hutter, 2017) decouples the $L_{2}$ regularization from the loss function and directly inject the weight decay, i.e., $\eta\lambda\bm{W}_{t}$, into the weight update. In this way, AdamW enforces stable shrinkage and tends to converge to smaller absolute values, particularly in high-magnitude regimes. Based on the update rule in Equation 1, we have
@@
Appendix D Limitations
@@
Appendix E Broader impact
@@
Appendix F Statement on the Use of Large Language Models
In the preparation of this paper, the authors used a Large Language Model solely for the purpose of improving the readability and language of the manuscript. Specifically, the LLM was used to polish and refine short passages of text originally written by the authors. All ideas, technical content, mathematical formulations, experimental design, data analysis, and conclusions are the sole work of the authors. The LLM was not used to generate any central ideas, creative content, or technical insights. The final manuscript was thoroughly reviewed and approved by all authors.

### v2

@@ 4.3 LLM Fine-Tuning
MOSS 241.8(+43.8%) 52.8 64.7 59.4
Table 4: Downstream evaluation accuracy (%) of Qwen-3-14B and Qwen-3-32B fine-tuned under BF16 and MOSS. All results are averaged over three independent runs.
Model Precision MATH500 GPQA-Diamond C-Eval AIME24 MMLU-Redux Qwen-3-14B BF16 96.8 62.1 86.2 79.1 88.6 MOSS 96.7 61.5 86.7 79.4 88.5 Qwen-3-32B BF16 98.3 64.2 88.2 79.3 89.6 MOSS 98.1 64.4 87.8 78.8 88.6
To further evaluate the scalability of MOSS on larger models, we fine-tune Qwen-3-14B and Qwen-3-32B under both BF16 and proposed MOSS framework, and report their downstream reasoning performance. Experiments are conducted on curated dataset of MegaScience fan2025megascience and NuminaMath-CoT li2024numinamath. All metrics are averaged over three independent fine-tuning runs to ensure statistical reliability. Table 4 presents the accuracy on five representative reasoning and knowledge benchmarks. Across both model sizes, MOSS matches the BF16 baseline on every task, demonstrating negligible performance loss despite operating at reduced precision. Notably, MOSS maintains strong stability on long-horizon reasoning tasks (e.g., MATH500, AIME24), where quantization methods often struggle with scale drift or cumulative error. These results indicate that MOSS scales effectively to 30B-class models while preserving competitive downstream performance.
4.4 Memory and Communication Gains
We conduct detailed profiling of peak activation memory usage and inter-GPU communication volume during LLaMA-2-7B fine-tuning on an 8×H200 cluster (3.2 TB/s aggregate NVLink bandwidth). All experiments use a per-GPU batch size of 4, a sequence length of 4096, and FSDP with ZeRO-2. Communication statistics are collected using the NCCL profiler and averaged over 1,000 training steps.
The measurements in Table 5 highlight the substantial memory-efficiency benefits introduced by MOSS. Even on bandwidth-rich H200 GPUs, MOSS delivers a 1.8× reduction in peak activation memory, confirming that the improvement arises from its FP8 microscaling algorithm rather than from hardware-level memory bandwidth limitations. We additionally evaluate compute–communication overlap using the PyTorch Profiler. MOSS achieves an overlap ratio exceeding 83%, indicating that it preserves strong scalability under high-speed NVLink interconnects and effectively reduces inter-GPU communication overhead during distributed training.
Table 5: Memory footprint and communication efficiency across BF16, COAT, and MOSS.
[…]
@@ 4.5 Ablation Study
4.5.1 GEMM Performance
To evaluate the efficiency of the GEMM kernel in MOSS, we conduct an ablation study comparing our implementation with several state-of-the-art FP8 quantization schemes, including Transformer Engine (TE)(TE), DeepGEMM from DeepSeek(deepgemm2025), and COAT (xi2024coat). This evaluation focuses on both raw GEMM performance and the practical trade-offs in quantization and hardware deployment.
Table 6: Runtime of quantized FP8 GEMM on NVIDIA H800 GPUs.
@@ 4.5.1 GEMM Performance
As shown in Table 6, MOSS achieves GEMM performance that is comparable to TE. Importantly, MOSS significantly outperforms COAT in GEMM throughput, highlighting the efficiency of our low-overhead quantization strategy. In comparison to DeepGEMM, our GEMM kernel runtime is relatively higher. This is expected, as DeepSeek leverages the Increasing Accumulation Precision technique to reduce dequantization overhead and employs extensive hardware-specific optimizations targeting NVIDIA’s Hopper GPUs. In contrast, our implementation is not tailored to Hopper GPUs, which lack the native Tensor Core support for MX data format. Despite this, MOSS remains competitive in performance across widely available hardware and maintains robust generalizability without relying on hardware-specific enhancements.
4.5.2 Quantization Fidelity
To evaluate the fidelity of different quantization strategies, we extracted activation tensors from LLaMA-2-7B during the actual fine-tuning process. Specifically, we sampled tensors from three critical layers—attention outputs, FFN intermediate activations, and LayerNorm inputs—every 100 training steps over a total of 10,000 steps. For each sample, we computed the signal-to-noise ratio (SNR) under three quantization schemes: 1) per-tensor quantization; 2) per-group quantization (group size = 128); 3) two-level microscaling (MOSS, micro group size = 32). The results, segmented into early ($\text{step}<2$K) and late ($\text{step}>8$K) training stages, are summarized in Table 7.
Table 7: SNR (dB) of activation tensors across different layers and quantization strategies during LLaMA-2-7B fine-tuning.
Layer Per-Tensor Per-Group MOSS (Two-Level)
Early Late Early Late Early Late
Attention Output 28.4 26.7 34.8 33.2 37.6 36.1
FFN Intermediate 25.9 24.1 32.3 30.7 35.9 35.3
LayerNorm Input 31.2 29.5 36.7 35.1 39.4 38.0
Geometric Mean 28.3 26.6 34.5 32.9 37.5 36.0
The experimental results confirm our theoretical analysis in Section 3.1. Specifically, MOSS consistently achieves superior quantization fidelity, improving SNR by 3.0–3.4 dB over per-group quantization and 9.2–9.4 dB over per-tensor quantization. More over, this SNR advantage is maintained across all layer types and throughout the training process, demonstrating the robustness of our two-level microscaling approach. These findings provide strong empirical support for the effectiveness of MOSS and validate the conclusions discussed in Section 3.1.
@@ 5.1 Low-Precision Training
Low-precision training refers to using low-precision bits, such as FP16/BF16, in the training process, as opposed to using FP32 full-precision throughout the training process. It is one of the most prominent directions in modern deep learning, as it offers reductions in computation costs, memory footprint and communication overhead (micikevicius2017mixed). Most existing training frameworks, e.g., Megatron-LLM (shoeybi2019megatron) and DeepSpeed (rasley2020deepspeed), adopt BF16 in LLM training by default. With the release of Nvidia Hopper GPUs, FP8 training (micikevicius2022fp8) becomes the next-generation low-precision technique. Transformer Engine (TE) from NVIDIA is the first framework designed for FP8 mixed-precision training, however, it only supports FP8 compute for linear layers while leaving all other operations in full precision. FP8-LM (peng2023fp8) extends FP8 quantization to gradients and first-order optimizer states, further improving the training throughput. However, they fail to reduce the memory footprint of activations, not fully exploiting the potential of memory and communication benefits of FP8 in fully sharded distributed training scenarios. COAT (xi2024coat) and DeepSeek-V3 (liu2024deepseek) both extend the per-group FP8 quantization to activations, achieving mixed granularity quantization. However, the fine-grained quantization strategy incurs additional computation cost in the dequantization process.
@@ 5.2 Scaling Techniques
Quantization transforms high-precision values within a given range into lower-precision values of a different range by tensor scaling. Scaling factor determination during training typically follows one of two strategies: Delayed scaling estimates the current scaling factor using a historical record of maximum absolute values from previous iterations. This approach assumes statistical consistency across iterations, making it vulnerable to outliers that violate this assumption and potentially destabilize training (fishman2024scaling). In contrast, just-in-time scaling dynamically adjusts scaling factors based on the current data distribution, typically using on-the-fly max-reduction. While more adaptive, this method introduces significant computational overhead—particularly in FP8 training—where multiple memory accesses can offset the performance gains of low-precision computation. Beyond these non-automatic scaling approaches, Unit Scaling (blake2023unit) recently introduced an automatic scaling method that maintains unit variance in weights and activations by applying static scaling constants throughout the network. However, Unit Scaling requires manual modification of the computational graph to insert unit-scaled operations and preserves certain “critical matmuls” (e.g., attention output projection and FFN down projection) in BF16 for stability. In contrast, MOSS offers a generic and automated scaling mechanism that integrates seamlessly into modern LLMs, achieving full FP8 precision without requiring precision fallbacks or manual graph modifications.
@@ Appendix C Discussion of Bound in AdamW Optimizer
Compared to Adam optimizer (kingma2014adam), AdamW (loshchilov2017decoupled) decouples the $L_{2}$ regularization from the loss function and directly inject the weight decay, i.e., $\eta\lambda\bm{W}_{t}$, into the weight update. In this way, AdamW enforces stable shrinkage and tends to converge to smaller absolute values, particularly in high-magnitude regimes. Based on the update rule in Equation 1, we have
@@
Appendix D Ablation Study on different intervals
To understand how the update interval affects the trade-off between computational overhead and model accuracy, we conduct an ablation study on LLaMA-2-7B fine-tuning with a range of interval values. Table 9 reports the corresponding scaling overhead, effective throughput, and downstream accuracy on NumGLUE.
Our findings reveal a clear pattern. Extremely frequent updates—such as the just-in-time configuration that recomputes scaling every step—introduce substantial overhead (3.8 ms/step) without improving throughput. In contrast, moderate update intervals in the range of 100–500 steps reduce scaling overhead to a negligible level (¡0.03 ms/step), yielding the highest effective throughput (1.043–1.044×) while matching or slightly exceeding BF16 accuracy. When the interval becomes too large (e.g., 2000), we observe a mild drop in accuracy due to insufficient tracking of scale drift, although the throughput remains comparable.
Overall, these results demonstrate that MOSS is highly robust to the choice of update interval across a broad regime. The default value of 500 steps strikes an optimal balance, minimizing scaling overhead while preserving accuracy and maximizing training efficiency.
Table 9: Impact of scaling interval on overhead, throughput, and accuracy during fine-tuning.
[…]
Appendix E End-to-End Throughput Improvement of Automatic Scaling
To quantify the end-to-end throughput benefits of different weight quantization strategies, we compared three scaling methods during LLaMA-2-7B fine-tuning on $\timesH8$00 GPUs:
  1. 1.
    Just-in-time (JIT) scaling: Performs on-the-fly max-reduction at every forward pass.
  2. 2.
    Delayed scaling: Uses historical maximum values with a moving window.
  3. 3.
    Automatic scaling (MOSS): Predicts scaling factors based on the learning rate schedule (our method).
Table 10: End-to-end throughput of different weight scaling strategies during LLaMA-2-7B fine-tuning on $\timesH8$00 GPUs.
Method
Scaling Overhead
Total Step
Time (ms)
Throughput
(token/s)
Speedup
[…]
Table 10 summarizes the end-to-end throughput during fine-tuning (batch size = 4, sequence length = 4096), averaged over 1,000 steps. As shown in Table 10, our automatic scaling method achieves an 8.7% end-to-end throughput improvement over the JIT baseline. This improvement primarily arises from eliminating the costly memory access pattern in JIT scaling, which requires reading the entire FP32 weight tensor from HBM for max-reduction. In contrast, MOSS incurs only negligible, constant-time overhead while maintaining accuracy.
Furthermore, to demonstrate the scalability and accuracy, we evaluated the downstream task accuracy of our automatic scaling method against just-in-time scaling under identical fine-tuning settings (LLaMA-2-7B on the MAmmoTH dataset). The results are summarized in Table 11.
Table 11: Downstream task accuracy comparison between JIT scaling and MOSS (Automatic). Differences are reported relative to JIT scaling.
Scaling Method Mathematics GSM8K NumGLUE
JIT Scaling 52.6% 64.9% 59.2%
MOSS (Automatic) 52.8% 64.7% 59.4%
Difference +0.2% -0.2% +0.2%
The observed differences in accuracy are within the statistical noise margin (±0.3% across three independent seeds), confirming that the theoretical upper bound in Equation (10) and the dynamic re-scaling mechanism provide sufficient numerical precision for practical training. Consequently, the 8.7% end-to-end throughput improvement achieved by automatic scaling is obtained without any degradation in model quality.
Appendix F Visualization of Quantization Methods
To give an overview about the trade-off between quantization fidelty and throughput, we visualize the performance landscape of different quantization strategies on a 2D plane (as shown in Figure 8), where the X-axis represents throughput (tokens/second) and the Y-axis represents quantization fidelity (SNR in dB). This visualization integrates our ablation study results and clearly illustrates that MOSS lies on the desirable Pareto frontier, effectively avoiding the core limitations exhibited by prior approaches.
Figure 8: Quantization Technique Trade-offs: Throughput vs. Precision.
Appendix G Limitations
@@
Appendix H Broader impact